# refrences.bib

### v1

@@ 3 Resolution of the feasible solution set
    Suppose that $e\in E$ and $\mathscr{I}_{j_{0}}(e)\neq\varnothing$ for some $j_{0}\in\mathscr{J}$. Also, without loss of generality, let $i_{0}=\max\mathscr{I}_{j_{0}}(e)$. So, according to Definition 6 , we have $\mathscr{I}_{j_{0}}(e)=\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\cup\left\{i_{0}\right\}$ and $S_{i_{0}j_{0}}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{j_{0}}\left(e,i_{0}\right)}S_{kj_{0}}^{\prime}\right)\neq\varnothing$. Thus, $\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\neq\varnothing$. To prove the converse statement, since $e(i)\in\mathscr{J}_{i}$ ($\forall i\in\mathscr{I}$), it is sufficient to show that if $e\left(i_{0}\right)=j_{0}$ and $\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\neq\varnothing$, then $S_{i_{0}j_{0}}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{j_{0}}\left(e,i_{0}\right)}S_{kj_{0}}^{\prime}\right)\neq\varnothing$. However, since $\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\subseteq\mathscr{I}_{j_{0}}(e)$, the statement $\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\neq\varnothing$ implies $\mathscr{I}_{j_{0}}(e)\neq\varnothing$. Hence, from the assumption, we have $\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\neq\varnothing$. Therefore, since $i_{0}\in\mathscr{I}_{j_{0}}(e)$ (because $e\left(i_{0}\right)=j_{0}$) and $\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\subseteq\mathscr{I}_{j_{0}}(e)$, it is concluded that $S_{i_{0}j_{0}}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{j_{0}}\left(e,i_{0}\right)}S_{kj_{0}}^{\prime}\right)\neq\varnothing$. ∎
Remark 2.
According to Corollary 3, the number of admissible functions is bounded above by $\prod_{i\in\mathscr{S}}\left|\mathscr{Z}_{i}\right|$, where $\left|\mathscr{Z}_{i}\right|$ denotes the cardinality of $\mathscr{Z}_{i}$. However, the actual number of admissible functions is usually much less than this value.
@@
    Let $x\in\bigcup_{e\in E}S(e)$. So, $x\in S\left(e_{0}\right)$ for some $e_{0}\in E$. Hence, according to (5), for each $j\in\mathscr{J}$ we have either $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}(e_{0})}S_{ij}^{\prime}$ (if $\mathscr{I}_{j}\left(e_{0}\right)\neq\varnothing$) or $x_{j}\in I_{j}$ (if $\mathscr{I}_{j}\left(e_{0}\right)=\varnothing$). But, since $S_{ij}^{\prime}=S_{ij}\cap I_{j}$ (Definition 4), from $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}\left(e_{0}\right)}S_{ij}^{\prime}$ it is obtained again $x_{j}\in I_{j}$. Consequently, $x_{j}\in I_{j}$, $\forall j\in\mathscr{J}$. On the other hand, from Definition 6 we have $e_{0}(i)=j_{i}\in\mathscr{J}_{i}\left(e_{0}\right)\subseteq\mathscr{J}_{i}$ ($\forall i\in\mathscr{I}$), which implies $\mathscr{I}_{j_{i}}\left(e_{0}\right)\neq\varnothing$, and therefore from (5) we have $x_{j_{i}}\in\bigcap_{k\in\mathscr{I}_{j_{i}}\left(e_{0}\right)}S_{kj_{i}}^{\prime}\subseteq S_{ij_{i}}^{\prime}$. Hence, $x_{j_{i}}\in S_{ij_{i}}^{\prime}$, $\forall i\in\mathscr{I}$. Now, Lemma 4 requires that $x\in S\left(A^{+},A^{-},b\right)$. Conversely, let $x\in S\left(A^{+},A^{-},b\right)$, $\mathscr{I}_{j}(x)=\left\{i\in\mathscr{I}:x_{j}\in S_{ij}^{\prime}\right\}$ and $\mathscr{J}_{i}(x)=\left\{j\in\mathscr{J}:x_{j}\in S_{ij}^{\prime}\right\}$. So, for each $i\in\mathscr{I}_{j}(x)$ and each $j\in\mathscr{J}_{i}(x)$, we have $S_{ij}^{\prime}\neq\varnothing$ that means $j\in\mathscr{J}_{i}$ (Definition 5). Also, Lemma 4 implies that $\mathscr{J}_{i}(x)\neq\varnothing$, $\forall i\in\mathscr{I}$. Without loss of generality, let $j_{i}=\min\mathscr{J}_{i}(x)$ and $e_{0}(i)=j_{i}$, $\forall i\in\mathscr{I}$. Therefore, $e_{0}$ is a function on $\mathscr{I}$ such that
@@
    Consequently, by Corollary 3 and (6) - (8) we have $e_{0}\in E$. Also, since $x\in S\left(A^{+},A^{-},b\right)$, then $x_{j}\in I_{j}$, $\forall j\in\mathscr{J}$ (Lemma 4). Particularly, $x_{j}\in I_{j}$, if $\mathscr{I}_{j}\left(e_{0}\right)=\varnothing$. This fact together with (5), (7) and (8) imply $x\in S\left(e_{0}\right)$. ∎
Suppose that $e\in E$, $x\in S(e)$ and $j\in\mathscr{J}$. So, According to (5), we have either $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}(e)}S_{ij}^{\prime}$ (if $\mathscr{I}_{j}(e)\neq\varnothing$) or $x_{j}\in I_{j}=\left[L_{j},U_{j}\right]$ (if $\mathscr{I}_{j}(e)=\varnothing$). On the other hand, in the former case, we have $\bigcap_{i\in\mathscr{I}_{j}(e)}S_{ij}^{\prime}\neq\varnothing$ (from Corollary 3). Consequently, by noting Remark 1, it follows that $\bigcap_{i\in\mathscr{I}_{j}(e)}S_{ij}^{\prime}$ is a closed and bounded (but, not necessarily connected). So, it is concluded that $S(e)$ ($\forall e\in E$) is indeed a closed bounded subset (and therefore, a compact subset) of $[0,1]^{n}$. This fact together with Theorem 1 imply that the feasible solutions set of Problem (1) is determined by the union of a finite number of compact sets, which are not necessarily connected.
Example 2.
Consider the problem stated in Example 1. According to Definition 5 and Table 4 (in Section 6), we obtain $\mathscr{J}_{1}=\{3,5\}$, $\mathscr{J}_{2}=\{1,2,7\}$, $\mathscr{J}_{3}=\{8,9\}$, $\mathscr{J}_{4}=\{2,3,4,7\}$, $\mathscr{J}_{5}=\{5\}$, $\mathscr{J}_{6}=\{8,9\}$ and $\mathscr{J}_{7}=\{2,6\}$. Hence, according to Remark 2, the number of admissible functions is bounded above by $\prod_{i\in\mathscr{I}}\left|\mathscr{J}_{i}\right|=2\times 3\times 2\times 4\times 1\times 2\times 2=192$. Now, noting Definition 6 and Corollary 3, consider functions $e_{1}=[3,2,9,2,5,9,6]$ and $e_{2}=[3,2,9,3,5,9,2]$ from $\mathscr{I}$ to $\bigcup_{i\in\mathscr{I}}\mathscr{J}_{i}$ so that $e_{p}(i)\in\mathscr{J}_{i}$ for any $i\in\mathscr{I}$ and $p\in\{1,2\}$. So, $e_{1}(2)=e_{1}(4)=2$ and $\mathscr{I}_{2}\left(e_{1},4\right)=\{2\}$ (Definition 6 (II)). Also, $S_{22}^{\prime}=\{0.75\}$ and $S_{42}^{\prime}=\{0.9\}$ (see Table 4). Therefore, $\mathscr{I}_{2}\left(e_{1},4\right)\neq\varnothing$ and $S_{42}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{2}\left(e_{1},4\right)}S_{k2}^{\prime}\right)=S_{42}^{\prime}\cap S_{22}^{\prime}=\varnothing$, which implies $2\notin\mathscr{J}_{4}\left(e_{1}\right)$ (Definition 6 ( III )). As a result, since $e_{1}(4)=2\notin\mathscr{J}_{4}\left(e_{1}\right)$, based on Definition 6 it follows that $e_{1}$ is not admissible. However, $e_{2}$ is indeed an admissible function. More precisely, we have $\mathscr{I}_{2}\left(e_{2},2\right)=\mathscr{I}_{9}\left(e_{2},3\right)=\mathscr{I}_{4}\left(e_{2},4\right)=\mathscr{I}_{5}\left(e_{2},5\right)=\mathscr{I}_{8}\left(e_{2},6\right)=\varnothing$ and $\mathscr{I}_{2}\left(e_{2},7\right)=\{2\}$. Also, for the non-empty set $\mathscr{I}_{2}\left(e_{2},7\right)$, we have $S_{72}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{2}(e,7)}S_{k2}^{\prime}\right)=S_{72}^{\prime}\cap S_{22}^{\prime}=\{0.75\}\neq\varnothing$. So, the three conditions of Definition 6 hold for $e_{2}$. Moreover, by noting Table 4, it follows that $\bigcap_{i\in\mathscr{I}_{1}\left(e_{2}\right)}S_{i1}^{\prime}=\bigcap_{i\in\mathscr{I}_{6}\left(e_{2}\right)}S_{i6}^{\prime}=\bigcap_{i\in\mathscr{I}_{7}\left(e_{2}\right)}S_{i7}^{\prime}=\varnothing$, $\bigcap_{i\in\mathscr{I}_{2}\left(e_{2}\right)}S_{i2}^{\prime}=S_{22}^{\prime}\cap S_{72}^{\prime}=\{0.75\}$, $\bigcap_{i\in\mathscr{I}_{3}\left(e_{2}\right)}S_{i3}^{\prime}=S_{13}^{\prime}=[0.1,0.3]\cup\{0.7\}$, $\bigcap_{i\in\mathscr{I}_{4}\left(e_{2}\right)}S_{i4}^{\prime}=S_{44}^{\prime}=[0,0.1]\cup[0.9,1]$, $\bigcap_{i\in\mathscr{I}_{5}\left(e_{2}\right)}S_{i5}^{\prime}=S_{55}^{\prime}=\{0.75\}$, $\bigcap_{i\in\mathscr{I}_{8}\left(e_{2}\right)}S_{i8}^{\prime}=S_{68}^{\prime}=[0.5,1]$ and $\bigcap_{i\in\mathscr{I}_{9}\left(e_{2}\right)}S_{i9}^{\prime}=S_{39}^{\prime}=\{0.2\}$. Furthermore, from Table 3, $I_{1}=[0,0.25]$, $I_{6}=[0.4,0.6]$ and $I_{7}=\{0.1\}$. Hence, from Definition $7,S\left(e_{2}\right)$ is calculated as the Cartesian product $S\left(e_{2}\right)=[0,0.25]\times\{0.75\}\times[0.1,0.3]\cup\{0.7\}\times[0,0.1]\cup[0.9,1]\times\{0.75\}\times[0.4,0.6]\times\{0.1\}\times[0.5,1]\times\{0.2\}$.
@@ 4 Simplification rules
3Compute sets $I_{ij}$, $S_{ij}$, $I_{j}$ and $S_{ij}^{\prime}$ for each $i\in\mathscr{I}$ and each $j\in\mathscr{J}$ (Corollary 1, Definition 4 and Table A2).
@@
5       stop; the problem is infeasible (Lemma 3 (a)).
@@
9             the problem is infeasible (Lemma 3 (b)).
@@
12             Apply the simplification rules 1 - 5 to determine the values of as many variables as possible. Delete the corresponding columns and redundant equations from the problem.
13            Find the set $S(e)$ for each admissible function $e\in E$ in the remaining problem (Definitions 6 and 7). Add the assigned variables (obtained in Step 4) to $S(e)$.
14            Generate the feasible solution set $S\left(A^{+},A^{-},b\right)$ by $\bigcup_{e\in E}S(e)$ (Theorem 1).
@@
It is to be noted that the FRE problems whose feasible regions are defined by the constraints $A\varphi x=b$ (i.e., $\max_{j=1}^{n}\left\{\varphi\left(a_{ij},x_{j}\right)\right\}=b_{i}$, $\forall i\in\mathscr{I}$), are indeed special cases of Problem (1). To see this, let $A^{+}=A$ and $A^{-}=\mathbf{0}_{m\times n}$ where $\mathbf{0}_{m\times n}$ is the zero matrix. So, $a_{ij}^{+}=a_{ij}$ and $a_{ij}^{-}=0$ ($\forall i\in\mathscr{I}$ and $\forall j\in\mathscr{J}$), and therefore each constraint $\max_{j=1}^{n}\left\{\max\left\{\varphi\left(a_{ij}^{+},x_{j}\right),\varphi\left(a_{ij}^{-},1-x_{j}\right)\right\}\right\}=b_{i}$ is reduced to $\max_{j=1}^{n}\left\{\max\left\{\varphi\left(a_{ij},x_{j}\right),\varphi\left(0,1-x_{j}\right)\right\}\right\}=b_{i}$. Furthermore, by the identity law of $\varphi$, it follows that $\varphi\left(0,1-x_{j}\right)=0$, $\forall j\in J$. Consequently, the above constraints are converted into $\max_{j=1}^{n}\left\{\varphi\left(a_{ij},x_{j}\right)\right\}=b_{i}$, $\forall i\in\mathscr{I}$. As a result, Algorithm 1 can be also used to solve the FRE problems. However, although the above algorithm accelerates the resolution of the problem by taking advantage of the five simplification rules, it should be noted that the number of $S(e)$ ($e\in E$) may grow exponentially in terms of the problem size. Consequently, finding other simplification rules or providing some methods (such as appropriate branch-and-bound techniques) is always in demand.
Example 3.
Let us look at the problem stated in Examples 1 and 2. The steps of Algorithm 1 are as follows:
@@
Tables 1 and 2 show all the sets $I_{ij}$ and $S_{ij}$ for each $i\in\mathscr{I}$ and $j\in\mathscr{J}$, respectively. In both tables, row $i$ ($i\in\mathscr{I}$) corresponds to equation $i$ and column $j$ ($j\in\mathscr{J}$) corresponds to variable $x_{j}$. By Table 1 and Definition 4, all the sets $I_{j}$ are shown in Table 3 for each $j\in\mathscr{J}$. Also, all the sets $S_{ij}^{\prime}$ ($\forall i\in\mathscr{I}$ and $\forall j\in\mathscr{J}$) are summarized in Table 4 by using Table 2, Table 3 and Definition 4.
@@
$[0,0.25]$$\{0.75\}$varnothing $\varnothing$ $\varnothing$ $\varnothing$ $\{0.1\}$varnothing $\varnothing$
@@
$[0,0.1]\cup[0.9,1]$
$\varnothing$ $\varnothing$ $\{0.1\}$varnothing $\varnothing$
$\varnothing$ $\{0.6\}$varnothing $\varnothing$ $\{0.75\}$varnothing $\varnothing$ $\varnothing$ $\varnothing$
@@
$\varnothing$ $[0.6,1]$varnothing $\varnothing$ $\varnothing$
@@
$[0.1,0.3]\cup\{0.7\}$
$\varnothing$ $[0.75,1]$varnothing $\varnothing$ $\varnothing$ $\varnothing$
$[0,0.25]$$\{0.75\}$varnothing $\varnothing$ $\varnothing$ $\varnothing$ $\{0.1\}$varnothing $\varnothing$
@@
$[0,0.1]\cup[0.9,1]$
$\varnothing$ $\varnothing$ $\{0.1\}$varnothing $\varnothing$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0.75\}$varnothing $\varnothing$ $\varnothing$ $\varnothing$
@@
$\varnothing$ $[0.75,0.9]$varnothing $\varnothing$ $\varnothing$
@@
Since $I_{7}=\{0.1\}$ is a singleton set (see Table 3) and $0.1\in S_{27}^{\prime}\cap S_{47}^{\prime}$ (see Table 4), Rule 2 indicates that $x_{7}=0.1$ for each feasible solution $x$ (particularly, $x_{7}^{*}=0.1$ for each optimal solution $x^{*}$). Also, column 7 and both rows 2 and 4 can be removed from the problem. So, by this simplification rule, the upper bound of the number of admissible functions is reduced from 192 (Example 2) to $\prod_{i\in\mathscr{I}-(2,4\}}\left|\mathscr{J}_{i}\right|=2\times 2\times 1\times 2\times 2=16$; that is $|E|\leq 16$. Subsequently, according to Table 4, it turns out that $\mathscr{J}_{5}=\{5\}$ and $S_{55}^{\prime}=\{0.75\}$ are singleton sets and $0.75\in S_{15}^{\prime}$. Thus, $x_{5}$ is assigned to $x_{5}=0.75$, and also column 5 and rows 1 and 5 can be deleted by Rule 4 . Consequently, we have $\prod_{i\in\mathscr{I}-(1,2,4,5)}\left|\mathscr{J}_{i}\right|=2\times 2\times 2=8$ and $|E|\leq 8$. It is to be noted that the first row can be also deleted by Rule 3; because, $S_{5j}^{\prime}\subseteq S_{1j}^{\prime}$, $\forall j\in\mathscr{J}$. Moreover, as shown in Tables 3 and 4, $S_{72}^{\prime}=I_{2}$. So, it is concluded that the seventh equation is also a redundant constraint and it can be deleted by Rule 5. Hence, by applying this simplification technique, the upper bound of the number of admissible functions is further reduced from $8$ to $|E|=\prod_{i\in\mathscr{I}-\{1,2,4,5,7\}}\left|\mathscr{J}_{i}\right|=2\times 2=4$. So, after applying the above simplification rules, columns $\{5,7\}$ and rows $\{1,2,4,5,7\}$ are deleted and we obtain $x_{5}^{*}=0.75$ and $x_{7}^{*}=0.1$. Therefore, the reduced matrices $A^{+}$and $A^{-}$, and the right-hand-side vector $b$ become
@@
The current matrices $A^{+}$and $A^{-}$are equivalent to two rows (rows 3 and 6 in the main problem) and seven columns (columns $1-4,6,8$ and 9 in the main problem). Furthermore, Table 4 is updated as follows:
Table 5: Updated Table 4 after applying Rules 1 - 5.
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $[0,0.2]\cup[0.8,1]$$\{0.2\}$
@@
Based on the results obtained in the previous step, we have $|E|=4$. Four admissible functions $e\in E$ are obtained from Table 5 and Definition 6 as $e_{1}=[6,6]$, $e_{2}=[6,7]$, $e_{3}=[7,6]$ and $e_{4}=[7,7]$. So, by Definition 7, their corresponding sets $S(e)$ are as follows: $S\left(e_{1}\right)=[0,1]\times[0,1]\times[0,1]\times[0,1]\times[0,1]\times[0.8,1]\times[0,1]$

### v2

@@ 3 Resolution of the feasible solution set
Suppose that $e\in E$ and $\mathscr{I}_{j_{0}}(e)\neq\varnothing$ for some $j_{0}\in\mathscr{J}$. Also, without loss of generality, let $i_{0}=\max\mathscr{I}_{j_{0}}(e)$. So, according to Definition 3.5 , we have $\mathscr{I}_{j_{0}}(e)=\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\cup\left\{i_{0}\right\}$ and $S_{i_{0}j_{0}}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{j_{0}}\left(e,i_{0}\right)}S_{kj_{0}}^{\prime}\right)\neq\varnothing$. Thus, $\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\neq\varnothing$. To prove the converse statement, since $e(i)\in\mathscr{J}_{i}$ ($\forall i\in\mathscr{I}$), it is sufficient to show that if $e\left(i_{0}\right)=j_{0}$ and $\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\neq\varnothing$, then $S_{i_{0}j_{0}}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{j_{0}}\left(e,i_{0}\right)}S_{kj_{0}}^{\prime}\right)\neq\varnothing$. However, since $\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\subseteq\mathscr{I}_{j_{0}}(e)$, the statement $\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\neq\varnothing$ implies $\mathscr{I}_{j_{0}}(e)\neq\varnothing$. Hence, from the assumption, we have $\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\neq\varnothing$. Therefore, since $i_{0}\in\mathscr{I}_{j_{0}}(e)$ (because $e\left(i_{0}\right)=j_{0}$) and $\mathscr{I}_{j_{0}}\left(e,i_{0}\right)\subseteq\mathscr{I}_{j_{0}}(e)$, it is concluded that $S_{i_{0}j_{0}}^{\prime}\cap\left(\bigcap_{k\in\mathscr{I}_{j_{0}}\left(e,i_{0}\right)}S_{kj_{0}}^{\prime}\right)\neq\varnothing$. ∎
Remark 3.8.
According to Corollary 3.7, the number of admissible functions is bounded above by $\prod_{i\in\mathscr{S}}\left|\mathscr{J}_{i}\right|$, where $\left|\mathscr{J}_{i}\right|$ denotes the cardinality of $\mathscr{J}_{i}$. However, the actual number of admissible functions is usually much less than this value.
@@
Let $x\in\bigcup_{e\in E}S(e)$. So, $x\in S\left(e_{0}\right)$ for some $e_{0}\in E$. Hence, according to (5), for each $j\in\mathscr{J}$ we have either $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}(e_{0})}S_{ij}^{\prime}$ (if $\mathscr{I}_{j}\left(e_{0}\right)\neq\varnothing$) or $x_{j}\in I_{j}$ (if $\mathscr{I}_{j}\left(e_{0}\right)=\varnothing$). But, since $S_{ij}^{\prime}=S_{ij}\cap I_{j}$ (Definition 2.7), from $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}\left(e_{0}\right)}S_{ij}^{\prime}$ it is obtained again $x_{j}\in I_{j}$. Consequently, $x_{j}\in I_{j}$, $\forall j\in\mathscr{J}$. On the other hand, from Definition 3.5 we have $e_{0}(i)=j_{i}\in\mathscr{J}_{i}\left(e_{0}\right)\subseteq\mathscr{J}_{i}$ ($\forall i\in\mathscr{I}$), which implies $\mathscr{I}_{j_{i}}\left(e_{0}\right)\neq\varnothing$, and therefore from (5) we have $x_{j_{i}}\in\bigcap_{k\in\mathscr{I}_{j_{i}}\left(e_{0}\right)}S_{kj_{i}}^{\prime}\subseteq S_{ij_{i}}^{\prime}$. Hence, $x_{j_{i}}\in S_{ij_{i}}^{\prime}$, $\forall i\in\mathscr{I}$. Now, Lemma 3.2 requires that $x\in S\left(A^{+},A^{-},b\right)$. Conversely, let $x\in S\left(A^{+},A^{-},b\right)$, $\mathscr{I}_{j}(x)=\left\{i\in\mathscr{I}:x_{j}\in S_{ij}^{\prime}\right\}$ and $\mathscr{J}_{i}(x)=\left\{j\in\mathscr{J}:x_{j}\in S_{ij}^{\prime}\right\}$. So, for each $i\in\mathscr{I}_{j}(x)$ and each $j\in\mathscr{J}_{i}(x)$, we have $S_{ij}^{\prime}\neq\varnothing$ that means $j\in\mathscr{J}_{i}$ (Definition 3.4). Also, Lemma 3.2 implies that $\mathscr{J}_{i}(x)\neq\varnothing$, $\forall i\in\mathscr{I}$. Without loss of generality, let $j_{i}=\min\mathscr{J}_{i}(x)$ and $e_{0}(i)=j_{i}$, $\forall i\in\mathscr{I}$. Therefore, $e_{0}$ is a function on $\mathscr{I}$ such that
@@
Consequently, by Corollary 3.7 and (6) - (8) we have $e_{0}\in E$. Also, since $x\in S\left(A^{+},A^{-},b\right)$, then $x_{j}\in I_{j}$, $\forall j\in\mathscr{J}$ (Lemma 3.2). Particularly, $x_{j}\in I_{j}$, if $\mathscr{I}_{j}\left(e_{0}\right)=\varnothing$. This fact together with (5), (7) and (8) imply $x\in S\left(e_{0}\right)$. ∎
Suppose that $e\in E$, $x\in S(e)$ and $j\in\mathscr{J}$. So, According to (5), we have either $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}(e)}S_{ij}^{\prime}$ (if $\mathscr{I}_{j}(e)\neq\varnothing$) or $x_{j}\in I_{j}=\left[L_{j},U_{j}\right]$ (if $\mathscr{I}_{j}(e)=\varnothing$). On the other hand, in the former case, we have $\bigcap_{i\in\mathscr{I}_{j}(e)}S_{ij}^{\prime}\neq\varnothing$ (from Corollary 3.7). Consequently, by noting Remark 1, it follows that $\bigcap_{i\in\mathscr{I}_{j}(e)}S_{ij}^{\prime}$ is a closed and bounded (but, not necessarily connected). So, it is concluded that $S(e)$ ($\forall e\in E$) is indeed a closed bounded subset (and therefore, a compact subset) of $0,1]^{n}$. This fact together with Theorem 1 imply that the feasible solutions set of Problem (1) is determined by the union of a finite number of compact sets, which are not necessarily connected.
Example 3.10.
[…]
@@ 4 Simplification rules
3Compute sets $I_{ij}$, $S_{ij}$, $I_{j}$ and $S_{ij}^{\prime}$ for each $i\in\mathscr{I}$ and each $j\in\mathscr{J}$ (Corollary 2.6, Definition 2.7 and Table A2).
@@
5   stop; the problem is infeasible (Lemma 3.1 (a)).
@@
9      the problem is infeasible (Lemma 3.1 (b)).
@@
12      Apply the simplification rules 4.1 - 4.5 to determine the values of as many variables as possible. Delete the corresponding columns and redundant equations from the problem.
13     Find the set $S(e)$ for each admissible function $e\in E$ in the remaining problem (Definitions 3.5 and 3.6). Add the assigned variables (obtained in Step 4) to $S(e)$.
14     Generate the feasible solution set $S\left(A^{+},A^{-},b\right)$ by $\bigcup_{e\in E}S(e)$ (Theorem 3.9).
@@
It is to be noted that the FRE problems whose feasible regions are defined by the constraints $A\varphi x=b$ (i.e., $\max_{j=1}^{n}\{\varphi(a_{ij}$, $x_{j})\}=b_{i}$, $\forall i\in\mathscr{I}$), are indeed special cases of Problem (1). To see this, let $A^{+}=A$ and $A^{-}=\mathbf{0}_{m\times n}$ where $\mathbf{0}_{m\times n}$ is the zero matrix. So, $a_{ij}^{+}=a_{ij}$ and $a_{ij}^{-}=0$ ($\forall i\in\mathscr{I}$ and $\forall j\in\mathscr{J}$), and therefore each constraint $\max_{j=1}^{n}\left\{\max\left\{\varphi\left(a_{ij}^{+},x_{j}\right),\varphi\left(a_{ij}^{-},1-x_{j}\right)\right\}\right\}=b_{i}$ is reduced to $\max_{j=1}^{n}\left\{\max\left\{\varphi\left(a_{ij},x_{j}\right),\varphi\left(0,1-x_{j}\right)\right\}\right\}=b_{i}$. Furthermore, by the identity law of $\varphi$, it follows that $\varphi\left(0,1-x_{j}\right)=0$, $\forall j\in J$. Consequently, the above constraints are converted into $\max_{j=1}^{n}\left\{\varphi\left(a_{ij},x_{j}\right)\right\}=b_{i}$, $\forall i\in\mathscr{I}$. As a result, Algorithm 1 can be also used to solve the FRE problems. However, although the above algorithm accelerates the resolution of the problem by taking advantage of the five simplification rules, it should be noted that the number of $S(e)$ ($e\in E$) may grow exponentially in terms of the problem size. Consequently, finding other simplification rules or providing some methods (such as appropriate branch-and-bound techniques) is always in demand.
Example 4.6.
Let us look at the problem stated in Examples 2.9 and 3.10. The steps of Algorithm 1 are as follows:
@@
Tables 1 and 2 show all the sets $I_{ij}$ and $S_{ij}$ for each $i\in\mathscr{I}$ and $j\in\mathscr{J}$, respectively. In both tables, row $i$ ($i\in\mathscr{I}$) corresponds to equation $i$ and column $j$ ($j\in\mathscr{J}$) corresponds to variable $x_{j}$. By Table 1 and Definition 2.7, all the sets $I_{j}$ are shown in Table 3 for each $j\in\mathscr{J}$. Also, all the sets $S_{ij}^{\prime}$ ($\forall i\in\mathscr{I}$ and $\forall j\in\mathscr{J}$) are summarized in Table 4 by using Table 2, Table 3 and Definition 2.7.
@@
$[0,0.25]$ $\{0.75\}$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0.1\}$ $\varnothing$ $\varnothing$
@@
$0,0.1]\cup[0.9,1]$
$\varnothing$ $\varnothing$ $\{0.1\}$ $\varnothing$ $\varnothing$
$\varnothing$ $\{0.6\}$ $\varnothing$ $\varnothing$ $\{0.75\}$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$
@@
$\varnothing$ $[0.6,1]$ $\varnothing$ $\varnothing$ $\varnothing$
@@
$0.1,0.3]\cup\{0.7\}$
$\varnothing$ $[0.75,1]$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$
$[0,0.25]$ $\{0.75\}$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0.1\}$ $\varnothing$ $\varnothing$
@@
$0,0.1]\cup[0.9,1]$
$\varnothing$ $\varnothing$ $\{0.1\}$ $\varnothing$ $\varnothing$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0.75\}$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$
@@
$\varnothing$ $[0.75,0.9]$ $\varnothing$ $\varnothing$ $\varnothing$
@@
Since $I_{7}=\{0.1\}$ is a singleton set (see Table 3) and $0.1\in S_{27}^{\prime}\cap S_{47}^{\prime}$ (see Table 4), Rule 4.2 indicates that $x_{7}=0.1$ for each feasible solution $x$ (particularly, $x_{7}^{*}=0.1$ for each optimal solution $x^{*}$). Also, column 7 and both rows 2 and 4 can be removed from the problem. So, by this simplification rule, the upper bound of the number of admissible functions is reduced from 192 (Example 3.10) to $\prod_{i\in\mathscr{I}-(2,4\}}\left|\mathscr{J}_{i}\right|=2\times 2\times 1\times 2\times 2=16$; that is $|E|\leq 16$. Subsequently, according to Table 4, it turns out that $\mathscr{J}_{5}=\{5\}$ and $S_{55}^{\prime}=\{0.75\}$ are singleton sets and $0.75\in S_{15}^{\prime}$. Thus, $x_{5}$ is assigned to $x_{5}=0.75$, and also column 5 and rows 1 and 5 can be deleted by Rule 4.4. Consequently, we have $\prod_{i\in\mathscr{I}-(1,2,4,5)}\left|\mathscr{J}_{i}\right|=2\times 2\times 2=8$ and $|E|\leq 8$. It is to be noted that the first row can be also deleted by Rule 4.3; because, $S_{5j}^{\prime}\subseteq S_{1j}^{\prime}$, $\forall j\in\mathscr{J}$. Moreover, as shown in Tables 3 and 4, $S_{72}^{\prime}=I_{2}$. So, it is concluded that the seventh equation is also a redundant constraint and it can be deleted by Rule 4.5. Hence, by applying this simplification technique, the upper bound of the number of admissible functions is further reduced from $8$ to $|E|=\prod_{i\in\mathscr{I}-\{1,2,4,5,7\}}\left|\mathscr{J}_{i}\right|=2\times 2=4$. So, after applying the above simplification rules, columns $\{5,7\}$ and rows $\{1,2,4,5,7\}$ are deleted and we obtain $x_{5}^{*}=0.75$ and $x_{7}^{*}=0.1$. Therefore, the reduced matrices $A^{+}$and $A^{-}$, and the right-hand-side vector $b$ become
@@
The current matrices $A^{+}$and $A^{-}$are equivalent to two rows (rows 3 and 6 in the main problem) and seven columns (columns 14,6,81-4,6,8 and 9 in the main problem). Furthermore, Table 4 is updated as follows:
Table 5: Updated Table 4 after applying Rules 4.1 - 4.5.
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $0,0.2]\cup[0.8,1]$ $\{0.2\}$
@@
Based on the results obtained in the previous step, we have $|E|=4$. Four admissible functions $e\in E$ are obtained from Table 5 and Definition 3.5 as $e_{1}=[6,6]$, $e_{2}=[6,7]$, $e_{3}=[7,6]$ and $e_{4}=[7,7]$. So, by Definition 3.6, their corresponding sets $S(e)$ are as follows: $S\left(e_{1}\right)=[0,1]\times[0,1]\times[0,1]\times[0,1]\times[0,1]\times[0.8,1]\times[0,1]$